OpenStudy (anonymous):

how to simplify log(10^5 x sqrt(x+1) / (x^2 - 4) x (x^2 +1)^6)

OpenStudy (anonymous):

\[\log(\frac{ 10^5 \times \sqrt{x+1} }{ (x^2 -4) \times (x^2 +1)^6}\]

OpenStudy (anonymous):

would this equal: \[5\log10 + \log \sqrt{x+1} - \log(x^2-4) - 6\log(x^2+1)\]